# Classification of abelian Schur groups I

### v1

@@ Abstract.
A finite group $G$ is called a Schur group if every $S$-ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of $\operatorname{Sym}(G)$ that contains all right translations of $G$. The list of all possible abelian Schur groups was obtained by Evdokimov, Kovács, and Ponomarenko in 2016. In two papers, we finish a classification of abelian Schur groups. In the present paper, we study schurity of several groups from the list. Namely, we prove that a direct product of the elementary abelian group of order 4 and a cyclic group, whose order is an odd prime power or a product of two distinct odd primes, is a Schur group and establish nonschurity of some other groups from the list.
@@ 1. Introduction
Let $G$ be a finite group. A subring of the group ring $\mathbb{Z}G$ is called an $S$-ring (Schur ring) over $G$ if it is a free $\mathbb{Z}$-module spanned by a partition of $G$ closed under taking inverse and containing the identity element $e$ of $G$ as a class. The notion of an $S$-ring was introduced by Schur [30]. Using the $S$-ring approach, Schur proved that every primitive permutation group containing a regular cyclic subgroup of composite order is $2$-transitive which generalizes the Burnside theorem [2, p. 339]. The general theory of $S$-rings was developed by Wielandt (see [32, Chapter IV]). To date, $S$-rings are realized as a powerful tool for studying permutation groups containing regular subgroups, combinatorial Cayley objects, especially, isomorphisms of them, representations of finite groups. For more details on $S$-rings and their applications, we refer the readers to [3, Section 2.4] and [20].
To study permutation groups having regular subgroups, Schur used $S$-rings whose partition of the underlying group arises from the action of an appropriate permutation group (see Section 2.3 for the exact definition). As it was shown by Wielandt [32], not all $S$-rings can be constructed in such way. An $S$-ring is said to be schurian if it arises from the action of a suitable permutation group [25]. The schurity property is one of the crucial properties of $S$-rings closely related to the isomorphism problem for Cayley graphs (see [20, Section 5.3] for details).
@@
In general, the above problem seems to be hard, in particular, because the number of $S$-rings over a given group is exponential in the order of the group. There are several results on schurity of nonabelian groups [21, 24, 26, 28, 29]. However, most of the results on the schurity problem are concerned with abelian groups. The first result on schurity of abelian groups was obtained in [25]. Namely, it was proved that cyclic $p$-groups of odd order are Schur. This result was extended to the case $p=2$ in [11]. In [16], a schurity of cyclic groups of order $pq$, where $p$ and $q$ are distinct primes, was proved. The above results were used later for solving the isomorphism problem for Cayley graphs over the corresponding groups [15, 16].
A complete classification of cyclic Schur groups was obtained in [9]. Namely, it was proved that a cyclic group of order $n$ is Schur if and only if $n$ belongs to one of the following families of integers:
@@
Investigations of schurity of abelian groups was continued in [10]. At first, necessary and sufficient condition of schurity for elementary abelian groups was obtained in this paper. Namely, an elementary abelian noncyclic group of order $n$ is Schur if and only if
@@
[10, Theorem 1.3] An abelian Schur group, which is neither cyclic nor elementary abelian, is isomorphic to a group from one of the following nine families:
@@
In the same paper, it was proved that the groups $E_{4}\times C_{p}$, where $p$ is an odd prime, are Schur groups. A schurity property of the groups $C_{2}\times C_{2^{k}}$ and $C_{3}\times C_{3^{k}}$ was established in [21] and [27], respectively. In [23], it was checked that $E_{9}\times C_{q}$ is a Schur group for every prime $q$.
@@
To prove Theorem 1.2, we characterize all $S$-rings over the groups from this theorem (see Theorem 7.1). We prove that every nontrivial $S$-ring over each of these groups arises from the action of an appropriate automorphism group or can be obtained as a tensor or generalized wreath product with special additional properties. To establish these properties is one of the crucial points in the proof because they allow to verify a sufficient condition of schurity for a generalized wreath product.
@@
The groups $C_{2p}\times C_{2^{k}}$, where $p$ is an odd prime and $k\geq 3$, are not Schur groups.
Since the class of Schur groups is closed under taking subgroups, it suffices to prove Theorem 1.3 only for $k=3$. Note that the groups $E_{4}\times C_{p}$ and $C_{2p}\times C_{4}$, where $p$ is an odd prime, are Schur groups. A schurity of the first group follows from [10, Theorem 1.4] as it was mentioned before, whereas a schurity of the second group will be verified in the subsequent paper together with a schurity of other abelian groups of order $8p$.
One more family of non-Schur groups from Theorem 1.1 is provided by the theorem below.
@@
In the proofs of Theorems 1.3 and 1.4, we construct new nonschurian $S$-rings over the groups $C_{2p}\times C_{8}$ and $E_{16}\times C_{p}$, respectively. These nonschurian $S$-rings are generalized wreath products of $S$-rings over proper subgroups of the above groups (see Sections 9 and 10). The nonschurian $S$-ring over $C_{2p}\times C_{8}$ from Section 9 in case $p=3$ was found by computer calculations using [14] and it is listed in [33]. The nonschurian $S$-ring over $E_{16}\times C_{p}$ from Section 10 in case $p=5$ was also found using [14].
We finish the introduction with a brief outline of the paper. Section 2 contains a necessary background of $S$-rings, especially, of the multiplier theorems, isomorphisms, and minimal $S$-rings. In Section 3, we provide an information on three constructions of $S$-rings, namely, tensor, generalized wreath, and star products. Several required facts on dual $S$-rings over the group of all irreducible complex characters of an abelian group are given in Section 4. Recent results on $S$-rings over an abelian group having a Sylow subgroup of prime order are provided in Section 5. In Section 6, we give a background of $S$-rings over cyclic groups, especially, over cyclic $p$-groups. In Section 7, we obtain a description of $S$-rings over the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$ (Theorem 7.1). In Sections 8, 9, and 10, we prove Theorems 1.21.3 and 1.4, respectively.
@@
If $f\in\operatorname{Sym}(\Omega)$ and $\Delta\subseteq\Omega$ is $f$-invariant, then the permutation of $\Delta$ induced by $f$ is denoted $f^{\Delta}$.
@@
The set of all positive divisors of $n\in\mathbb{Z}$ is denoted by $[n]$.
Given a cyclic group $G$ of order $n$ and $l\in[n]$, the subgroup of $G$ of order $l$ is denoted by $G_{l}$ and the set of all generators of $G_{l}$ is denoted by $G_{l}^{*}$.
If $X\subseteq H\times G$, where $G$ is a cyclic group of order $n$, $h\in H$, and $l\in[n]$, then put $X_{h}=h^{-1}X\cap G$ and $X_{h,l}=h^{-1}X\cap G_{l}^{*}$.
@@ 2.1. Definitions and basic facts
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, $\mathbb{Z}G$ is an $S$-ring. The partition $\{\{e\},G^{\#}\}$ defines the $S$-ring $\mathcal{T}_{G}$ of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$ and antisymmetric if for every $X\in\mathcal{S}(\mathcal{A})$ the equality $X=X^{-1}$ implies that $X=\{e\}$.
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that $\underline{X}\cdot\underline{Y}=\sum_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}$. Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$.
@@
[10, Lemma 2.1] Let $\mathcal{A}$ be an $S$-ring over a group $G$, $H$ an $\mathcal{A}$-subgroup of $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then the number $|X\cap Hx|$ does not depend on $x\in X$.
@@
[21, Theorem 2.6] Let $X$ be a basic set of an $S$-ring $\mathcal{A}$ over a group $G$. Suppose that $H\leq\operatorname{rad}(X\setminus H)$ for some subgroup $H$ of $G$ such that $X\cap H\neq\varnothing$ and $X\setminus H\neq\varnothing$. Then $\operatorname{rad}(X)\leq H$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$.
@@ 2.2. Multiplier theorems
Let $G$ be an abelian group. If $X,Y\subseteq G$ are such that $Y^{(m)}=X$ for some $m\in\mathbb{Z}$ coprime to $|G|$, then we say that $X$ and $Y$ are rationally conjugate. In this case, if $X\in\operatorname{Orb}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $Y\in\operatorname{Orb}(K,G)$. The following two statements are known as the first and second Schur theorems on multipliers (see [32, Theorem 23.9, (a)-(b)]).
@@
Let $G=H\times D$, where $H$ is an abelian group and $D\cong C_{n}$ with $n$ coprime to $|H|$, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then there is $K\leq\operatorname{Aut}(D)$ such that every nonempty set $X_{h,l}=h^{-1}X\cap D_{l}^{*}$, where $h\in H$ and $l\in[n]$, is an orbit of $K$.
@@
Let $K$ be the setwise stabilizer of $X$ in the group $M\leq\operatorname{Aut}(G)$ such that $M^{H}$ is trivial and $M^{D}=\operatorname{Aut}(D)$. By the definition of $K$, each set $X_{h,l}$ is $K$-invariant. On the other hand, for $x,y\in X_{h,l}$, there is an integer $m$ coprime to $n$ such that $x^{m}=y$ and $m\equiv 1\mod|H|$ and hence $(hx)^{m}=hy$. The latter implies that $X^{(m)}\cap X\neq\varnothing$. So $X^{(m)}=X$ by Lemma 2.3. Therefore $\sigma_{m}\in K$. Thus, $K$ is transitive on each $X_{h,l}$ and consequently $X_{h,l}$ is an orbit of $K$. ∎
@@ 2.3. Isomorphisms and schurity
Let $K$ be a subgroup of $\operatorname{Sym}(G)$ containing $G_{r}$. Schur proved in [30] that the $\mathbb{Z}$-submodule
@@
is an $S$-ring over $G$. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. The group $G$ is called a Schur group if every $S$-ring over $G$ is schurian.
If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=V(K^{S},S)$. This yields that $\mathcal{A}_{S}$ is also schurian. Therefore a section (in particular, a subgroup) of a Schur group is also a Schur group.
Clearly, $\operatorname{Aut}(\mathcal{A})\geq K$. One can verify that $\mathcal{A}$ is schurian if and only if
@@
or, equivalently, $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(\operatorname{Aut}(\mathcal{A})_{e},G)$. It is easy to see that $\mathcal{T}_{G}=V(K,G)$ for every $2$-transitive group $K\leq\operatorname{Sym}(G)$ containing $G_{r}$.
The $S$-ring $\mathcal{A}$ is said to be normal if $G_{r}$ is normal in $\operatorname{Aut}(\mathcal{A})$ or, equivalently, $\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)=G_{r}\rtimes\operatorname{Aut}(G)$.

### v3

@@ Abstract.
A finite group $G$ is called a Schur group if every Schur ring over $G$ is schurian, i.e. associated in a natural way with a subgroup of the symmetric group $\operatorname{Sym}(G)$ that contains all right translations of $G$. The list of all possible abelian Schur groups was obtained by Evdokimov, Kovács, and Ponomarenko in 2016. In two papers, we complete a classification of abelian Schur groups. In the present paper, we study schurity of several groups from the list. First, we prove that a direct product of the elementary abelian group of order 4 and a cyclic group, whose order is an odd prime power or a product of two distinct odd primes, is a Schur group. Second, we establish nonschurity of some other groups from the list.
@@ 1. Introduction
Let $G$ be a finite group. A subring of the group ring $\mathbb{Z}G$ is called an $S$-ring (a Schur ring) over $G$ if it is a free $\mathbb{Z}$-module spanned by a partition of $G$, closed under taking inverse and containing the identity element $e$ of $G$ as a class. The notion of $S$-ring was introduced by Schur [32]. Using the $S$-ring approach, Schur proved that every primitive permutation group containing a regular cyclic subgroup of composite order is $2$-transitive which generalizes the Burnside theorem [2, p. 339]. The general theory of $S$-rings was developed by Wielandt (see [34, Chapter IV]). To date, $S$-rings are realized as a powerful tool for studying permutation groups containing regular subgroups, combinatorial Cayley objects, especially, isomorphisms of them, and representations of finite groups. For more details on $S$-rings and their applications, we refer the reader to [3, Section 2.4] and [22].
To study permutation groups having regular subgroups, Schur used $S$-rings whose partition of the underlying group consists of the orbits of a one-point stabilizer of an appropriate permutation group (see Section 2.3 for the exact definition). As was shown by Wielandt [34], not all $S$-rings can be constructed in this way. An $S$-ring is said to be schurian if it can [27]. The schurity (the property to be schurian) is one of the crucial properties of $S$-rings, closely related to the isomorphism problem for Cayley graphs (see [22, Section 5.3] for details).
@@
In general, the above problem seems to be hard, in particular, because the number of $S$-rings over a given group can be exponential in the order of the group. There are several results on schurity of nonabelian groups [23, 26, 28, 30, 31]. However, most of the results on the schurity problem are concerned with abelian groups. The first result on schurity of abelian groups was obtained in paper [27]. Namely, it was proved that cyclic $p$-groups of odd order are Schur. This result was extended to the case $p=2$ in paper [12]. In paper [18], the schurity of cyclic groups of order $pq$, where $p$ and $q$ are distinct primes, was proved. The above results were used later for solving the isomorphism problem for Cayley graphs over the corresponding groups [17, 18].
A complete classification of cyclic Schur groups was obtained in paper [10]. Namely, it was proved that a cyclic group of order $n$ is Schur if and only if $n$ belongs to one of the following families of integers:
@@
Investigations of schurity of abelian groups was continued in paper [11]. At first, necessary and sufficient condition of schurity for elementary abelian groups was obtained there. Namely, an elementary abelian noncyclic group of order $n$ is Schur if and only if
@@
[11, Theorem 1.3] An abelian Schur group, which is neither cyclic nor elementary abelian, is isomorphic to a group from one of the following nine families:
@@
In the same paper, it was proved that the groups $E_{4}\times C_{p}$, where $p$ is an odd prime, are Schur groups. A schurity property of the groups $C_{2}\times C_{2^{k}}$ and $C_{3}\times C_{3^{k}}$ was established in papers [23] and [29], respectively. In paper [25], it was checked that $E_{9}\times C_{q}$ is a Schur group for every prime $q$.
@@
To prove Theorem 1.2, we characterize all $S$-rings over the groups from this theorem.
Theorem 1.3.
Every nontrivial $S$-ring over $E_{4}\times C_{p^{k}}$ or $E_{4}\times C_{pq}$, where $p$ and $q$ are odd primes and $k\geq 1$, is cyclotomic or a nontrivial tensor or generalized wreath product.
A similar statement was proved for an arbitrary cyclic group in papers [20, 21] and for some other abelian groups close to cyclic groups in papers [11, 22, 25, 29]. On the other hand, the above statement does not hold, e.g., for elementary abelian groups (see [34]). It seems interesting to characterize all groups whose all nontrivial $S$-rings are cyclotomic or nontrivial tensor or generalized wreath products (see Subsection 2.3 and Section 3 for the definitions).
@@
The group $C_{2p}\times C_{2^{k}}$, where $p$ is an odd prime and $k\geq 3$, is not a Schur group.
Since the class of Schur groups is closed under taking subgroups, it suffices to prove Theorem 1.4 only for $k=3$. It should be noted that if $k=1,2$, then the group from Theorem 1.4 is a Schur group. Indeed, the schurity of the group $E_{4}\times C_{p}$ (case $k=1$) follows from [11, Theorem 1.4] as was mentioned before, whereas the schurity of the group $C_{2p}\times C_{4}$ (case $k=2$) will be verified in the subsequent paper together with the schurity of other abelian groups of order $8p$, where $p$ is an odd prime.
One more family of non-Schur groups from Theorem 1.1 is presented by the theorem below.
@@
In the proofs of Theorems 1.4 and 1.5, we construct new nonschurian $S$-rings over the groups $C_{2p}\times C_{8}$ and $E_{16}\times C_{p}$, respectively. These nonschurian $S$-rings are generalized wreath products of $S$-rings over proper subgroups of the above groups (see Sections 9 and 10). The nonschurian $S$-ring over $C_{2p}\times C_{8}$ from Section 9 in case $p=3$ was found by computer calculations using the package [16] and appeared in paper [35].
We finish the introduction with a brief outline of the paper. Section 2 contains a necessary background of $S$-rings, especially, of the multiplier theorems and isomorphisms. In Section 3, we provide an information on three constructions of $S$-rings, namely, tensor, generalized wreath, and star products. Several required facts on dual $S$-rings over the group of all irreducible complex characters of an abelian group are given in Section 4. Recent results on $S$-rings over an abelian group having a Sylow subgroup of prime order are provided in Section 5. In Section 6, we give a background of $S$-rings over cyclic groups, especially, over cyclic $p$-groups. In Section 7, we obtain a description of $S$-rings over the groups $E_{4}\times C_{p^{k}}$ and $E_{4}\times C_{pq}$ (Theorem 7.1) that implies Theorem 1.3. In Sections 8, 9, and 10, we prove Theorems 1.21.4 and 1.5, respectively.
@@
If $f\in\operatorname{Sym}(\Omega)$ and $\Delta\subseteq\Omega$ is such that $\Delta^{f}=\Delta$, then the permutation of $\Delta$ induced by $f$ is denoted $f^{\Delta}$.
@@
The set of all positive divisors of $n\in\mathbb{Z}$ is denoted by $\operatorname{Div}(n)$.
Given a cyclic group $D$ of order $n$ and $l\in\operatorname{Div}(n)$, the subgroup of $D$ of order $l$ is denoted by $D_{l}$ and the set of all generators of $D_{l}$ is denoted by $D_{l}^{*}$.
If $X\subseteq H\times D$, where $D$ is a cyclic group of order $n$, $h\in H$, and $l\in\operatorname{Div}(n)$, then put $X_{h}=h^{-1}X\cap D$ and $X_{h,l}=h^{-1}X\cap D_{l}^{*}$.
@@ 2.1. Definitions and basic facts
The elements of $\mathcal{S}$ are called the basic sets of $\mathcal{A}$. The number of basic sets is called the rank of $\mathcal{A}$ and denoted by $\operatorname{rk}(\mathcal{A})$. Clearly, the group ring $\mathbb{Z}G$ is an $S$-ring. If $|G|\neq 1$, then the partition $\{\{e\},G^{\#}\}$ defines the $S$-ring $\mathcal{T}_{G}$ of rank $2$ over $G$. It is easy to check that if $X,Y\in\mathcal{S}(\mathcal{A})$, then $XY\in\mathcal{S}(\mathcal{A})$ whenever $|X|=1$ or $|Y|=1$. The $S$-ring $\mathcal{A}$ is called symmetric if $X=X^{-1}$ for every $X\in\mathcal{S}(\mathcal{A})$ and antisymmetric if for every $X\in\mathcal{S}(\mathcal{A})$ the equality $X=X^{-1}$ implies that $X=\{e\}$.
Let $X,Y\in\mathcal{S}$. If $Z\in\mathcal{S}$, then by Condition $(3)$, the number of distinct representations of $z\in Z$ in the form $z=xy$ with $x\in X$ and $y\in Y$ does not depend on the choice of $z\in Z$. Denote this number by $c^{Z}_{XY}$. One can see that
$\underline{X}\cdot\underline{Y}=\sum_{Z\in\mathcal{S}(\mathcal{A})}c^{Z}_{XY}\underline{Z}.$
Therefore the numbers $c^{Z}_{XY}$ are the structure constants of $\mathcal{A}$ with respect to the basis $\{\underline{X}:\ X\in\mathcal{S}\}$.
@@
[11, Lemma 2.1] Let $\mathcal{A}$ be an $S$-ring over a group $G$, $H$ an $\mathcal{A}$-subgroup of $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then the number $|X\cap Hx|$ does not depend on $x\in X$.
@@
[23, Theorem 2.6] Let $X$ be a basic set of an $S$-ring $\mathcal{A}$ over a group $G$. Suppose that $H\leq\operatorname{rad}(X\setminus H)$ for some subgroup $H$ of $G$ such that $X\cap H\neq\varnothing$ and $X\setminus H\neq\varnothing$. Then $\operatorname{rad}(X)\leq H$ and $X=\langle X\rangle\setminus\operatorname{rad}(X)$.
@@ 2.2. Multiplier theorems
Let $G$ be an abelian group. If $X,Y\subseteq G$ are such that $Y^{(m)}=X$ for some $m\in\mathbb{Z}$ coprime to $|G|$, then we say that $X$ and $Y$ are rationally conjugate. In this case, if $X\in\operatorname{Orb}(K,G)$ for some $K\leq\operatorname{Aut}(G)$, then $Y\in\operatorname{Orb}(K,G)$. The following two statements are known as the first and second Schur theorems on multipliers (see [34, Theorem 23.9, (a)-(b)]).
@@
Let $G=H\times D$, where $H$ is an abelian group and $D\cong C_{n}$ with $n$ coprime to $|H|$, $\mathcal{A}$ an $S$-ring over $G$, and $X\in\mathcal{S}(\mathcal{A})$. Then there exists $K^{D}\leq\operatorname{Aut}(D)$ satisfying the following: $X_{h,l}=h^{-1}X\cap D_{l}^{*}\in\operatorname{Orb}(K^{D},D)$ for all $h\in H$ and $l\in\operatorname{Div}(n)$ such that $X_{h,l}\neq\varnothing$.
@@
Since $|H|$ and $n$ are coprime, $\operatorname{Aut}(G)=\operatorname{Aut}(H)\times\operatorname{Aut}(D)$. Let $K$ be the setwise stabilizer of $X$ in the group $\{\operatorname{id}_{H}\}\times\operatorname{Aut}(D)\leq\operatorname{Aut}(G)$ and $h\in H$ and $l\in\operatorname{Div}(n)$ be such that $X_{h,l}\neq\varnothing$. By the definition of $K$, the set $X_{h,l}$ is $K$-invariant. On the other hand, for all $x,y\in X_{h,l}$, there is an integer $m$ coprime to $n$ such that $x^{\sigma_{m}}=x^{m}=y$ and $m\equiv 1\mod|H|$ and hence $(hx)^{m}=hy$. The latter implies that $X^{(m)}\cap X\neq\varnothing$. So $X^{(m)}=X$ by Lemma 2.3. Therefore $\sigma_{m}\in K$ and consequently $K$ is transitive on $X_{h,l}$. Thus, $X_{h,l}$ is an orbit of $K$ and hence of the group $K^{D}$ induced by $K$ on $D$. ∎
@@ 2.3. Isomorphisms and schurity
Let $K$ be a subgroup of $\operatorname{Sym}(G)$ containing $G_{r}$. Schur proved in paper [32] that the $\mathbb{Z}$-submodule
@@
is an $S$-ring over $G$. An $S$-ring $\mathcal{A}$ over $G$ is called schurian if $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ with $K\geq G_{r}$. One can see that $\mathcal{T}_{G}=V(K,G)$ for every $2$-transitive group $K\leq\operatorname{Sym}(G)$ containing $G_{r}$ and hence $\mathcal{T}_{G}$ is schurian. The group $G$ is called a Schur group if every $S$-ring over $G$ is schurian.
Let $\mathcal{A}=V(K,G)$ for some $K\leq\operatorname{Sym}(G)$ such that $K\geq G_{r}$. One can see that $\operatorname{Aut}(\mathcal{A})\geq K$. If $S$ is an $\mathcal{A}$-section, then $\mathcal{A}_{S}=V(K^{S},S)$. This yields that $\mathcal{A}_{S}$ is also schurian. Therefore a section (in particular, a subgroup) of a Schur group is also a Schur group.
One can verify that $\mathcal{A}$ is schurian if and only if
@@
or, equivalently, $\mathcal{S}(\mathcal{A})=\operatorname{Orb}(\operatorname{Aut}(\mathcal{A})_{e},G)$.
The $S$-ring $\mathcal{A}$ is said to be normal if $G_{r}$ is normal in $\operatorname{Aut}(\mathcal{A})$ or, equivalently,
$\operatorname{Aut}(\mathcal{A})\leq\operatorname{Hol}(G)=G_{r}\rtimes\operatorname{Aut}(G).$